A piece of Aluminium is weighed in air and has an apparent mass of 4.0000 kg (that many

significant digits). What is its true weight ?
Answer #1

Remember the Arcmedes Principle:

"a body totally or partially immersed in a fluid is subject to an upward force equal in magnitude to the weight of fluid it displaces."

Now using this principle:

mass of aluminium= 4.0000 kg

Density of aluminium = 2700.0 kg/m^3

Volume of aluminium = Mass/density

Volume = (4.0000/2700.0)*1000 = 1.4815 L

Now density of Air = 1.2041 gm/L

Volume displaced = 1.4815*1.2041 = 1.7839 gm of Air

Apparent loss of mass in aluminium = 1.7839 gm

True weight = 4.0000 kg + 1.7839 gm = 4.0019 kg